Solve the system of equations:
2x-3y+2z=-1
x+2y=14
x-3z=-5

Answer by Alan3354(21608)

2x-3y+2z=-1
x+2y=14 --> y = -x/2 + 7
x-3z=-5 --> z = (x+5)/3
------------
Sub for y and z in the 1st eqn
2x - 3(-x/2 + 7) + 2(x+5)/3 = -1
Multiply by 6
12x + 9x -126 + 4x + 20 = -6
25x = 100
x = 4
------
Sub for x into 2nd and 3rd eqns to find y and z.